OpenStudy (anonymous):

Can somebody explain to me how to graph Conic Sections and how to find the center, foci , and all of that from an equation and how to take for example "16x^2 -81y^2 + 324 y – 1620 = 0" into an equation we can graph... just everything.. I need help <3 My teacher isnt helping at all. ;-;

OpenStudy (anonymous):

So are you familiar with the equations of each type of conic section?

OpenStudy (anonymous):

... no....

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

hiya, can you help me.....? ;-;

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

You're the best. omg

OpenStudy (anonymous):

just let me know

OpenStudy (ybarrap):

Have you looked at this? http://en.wikipedia.org/wiki/Conic_section#Features

OpenStudy (anonymous):

I havent..

OpenStudy (ybarrap):

Look especially at this section - http://en.wikipedia.org/wiki/Conic_section#Cartesian_coordinates

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

So the equation up top is a circle? where does the y come from?

OpenStudy (anonymous):

324y*

OpenStudy (anonymous):

or is it a parabola ?

OpenStudy (ybarrap):

So we have $$ 16x^2 -81y^2 + 324 y – 1620 = 0 $$ $$ Ax^2 + Bxy + Cy^2 +Dx + Ey + F = 0 $$ What is A-F? Then use $$ B^2-4AC $$ To determine which type of conic section using the last link I sent. In other words, is the last equation, <,> or equal to zero?

OpenStudy (ybarrap):

Do you see where the "y" comes from now? For example, A=16. See that?

OpenStudy (ybarrap):

BRB, figure out the rest -- I'll check your work.

OpenStudy (anonymous):

Yea, But what about the parts that aren't there like Bzy and Dz?

OpenStudy (ybarrap):

Those would be zero then. BRB

OpenStudy (anonymous):

Okay... so B=0? A=16 and C=-81? so 0^2-4(16*-81) =5184 5184>0 so it'd be a hyperbola?

OpenStudy (anonymous):

In my assignment i need this for i need to graph the points and the equation. for example: Find the graphing form, coordinates of the center & foci, the length of the transverse and conjugate axis, then graph the points and equation. 16x2 -81y2 + 324 y – 1620 = 0

OpenStudy (ybarrap):

yes, hyperbola

OpenStudy (ybarrap):

So we now think its a hyperbola, what is the equation of a hyperbola? According to the 1st link I sent what is the equation - http://en.wikipedia.org/wiki/Conic_section#Features

OpenStudy (anonymous):

X^2/a^2-y^2/b^2=1

OpenStudy (ybarrap):

Yes, so to get "a" and "b", we'll need this - http://en.wikipedia.org/wiki/Hyperbola#Quadratic_equation

OpenStudy (anonymous):

there's so much work to this section of algebra, it's crazy..

OpenStudy (ybarrap):

It is, yes. But it is just a little tedious but not hard. So we were correct that this is a hyperbola and you have the recipe to find a and b. Here is the answer but don't just accept it. Check it using the recipe - http://www.wolframalpha.com/input/?i=16x%5E2+-81y%5E2+%2B+324+y+%E2%80%93+1620+%3D+0

OpenStudy (anonymous):

Why have i never found this website????????

OpenStudy (ybarrap):

Just use it to check your work - don't let it stop you from learning pls. $$ a^{2} = -\frac{\Delta}{\lambda_{1}D} = -\frac{\Delta}{\lambda_{1}^{2}\lambda_{2}}\\ b^{2} = -\frac{\Delta}{\lambda_{2}D} = -\frac{\Delta}{\lambda_{1}\lambda_{2}^{2}} $$ Where all are defined in the last link I sent.

OpenStudy (anonymous):

whats that sideways T looking thing mean?

OpenStudy (ybarrap):

Those are the two roots of the following equation: $$ \lambda^{2} - \left( A_{xx} + A_{yy} \right)\lambda + D = 0 $$

OpenStudy (ybarrap):

It's called "Lambda1" and "Lambda2." Lambda \(\Delta\) is a Greek symbol (equivalent to our letter "L")

OpenStudy (ybarrap):

When you solve this last equation, because it is a quadratic, it will have 2 roots.

OpenStudy (ybarrap):

Here is the meaning of \(A_{xx}\) and \( A_{yy}\) - $$ A_{xx} x^{2} + 2 A_{xy} xy + A_{yy} y^{2} + 2 B_{x} x + 2 B_{y} y + C = 0 $$ That's the same equation we had before but with A-F replaced by new names.

OpenStudy (ybarrap):

For example \( A_{yy}=-81\) from your equation. Make sense?

OpenStudy (ybarrap):

You're not giving up on me, are you?

OpenStudy (ybarrap):

Once you do this once, the second time will be easy! Just seems difficult the 1st time.

OpenStudy (ybarrap):

You were asking good questions.

OpenStudy (anonymous):

No sorry i looked off for a moment, was doing something and also reading what you were typing

OpenStudy (ybarrap):

I think you have everything you need.

OpenStudy (anonymous):

So far i understand how to find what type of conic section. but how about how to find the Verti, Foci, and Center? or the length of the transverse and conjugate axis?

OpenStudy (ybarrap):

Center is here http://en.wikipedia.org/wiki/Hyperbola#Quadratic_equation Using "Features" (i.e. e,c,l and p) here you can find the other parameters http://en.wikipedia.org/wiki/Conic_section#Features You will see how everything ties together here - http://en.wikipedia.org/wiki/Hyperbola#Nomenclature_and_features

OpenStudy (anonymous):

okay ill check it out thanks

OpenStudy (ybarrap):

send me note if after your review you have questions -- I know you can do this. Show me what you tried take care

OpenStudy (ybarrap):

Here's a little animation that might help -  http://assets.openstudy.com/updates/attachments/555fcf87e4b06a1264ab6fc1-ybarrap-1432483249200-hyperbola.gif

OpenStudy (ybarrap):

Here is where you can interact with my simulation - http://tube.geogebra.org/student/m1235825